OpenStudy (mtalhahassan2):

Need help!! 8. Solve for theta. 0

OpenStudy (campbell_st):

factor out cos \[\cos(\theta)(5\sin(\theta) -3) \] now set each factor to zero and solve \[\cos(tehta) = 0~~~and~~~5\sin(\theta) - 3 = 0\]

OpenStudy (mtalhahassan2):